OpenStudy (anonymous):

A 720ml container holds a mixture of red, yellow & white paint in a 5:3:12 ratio. How much more red paint than yellow is present in the mixture?

OpenStudy (turingtest):

5x+3x+12x=720 x=36 red paint=5(36)=180=r yellow paint=3(36)=108=y r-y=180=108=72ml

OpenStudy (turingtest):

^last line should be r-y=180-108=72ml